Online analysis: results and TODO

From Hall A Wiki
Jump to: navigation, search

Back to DVCS

This page is in progress, please new items to it. When you obtain a result, log in into the Halog and link it to this page. if you are willing to take charge of a given analysis, please indicate it on this page as well as the date at which you are starting to work on it.

Matching data taking with run plans

  • Update the how to
  • Maintain db_run.dat file: Make sure that every LHRS momentum and angle changes are logged (Florian Itard is keeping up with that)
  • Database of run indicating line by line the run number, the target in position, the average beam current, LHRs pnom, LHRS angle, the purpose of the run

Luminosity, efficiency

Boiling test

Initial calibration on 10-17 owl shift: [shift summary][initial plan]
Second calibration on 10-21 swing shift: [run list]

Trigger efficiency

Online analysis by B. Michaels [here]. 

Run7711 T3 and T4 Triggers

S0 hit pattern vs x and y (at vdc) shows efficiency in excess of 99% except at extreme ends of focal plane halog entry 336817

Cut script saved at /adaqfs/home/adaq/dvcs/onlana/GoodHRS.h TCut cGoodVDC = "<1."; TCut cGoodAngles = "abs(<0.1&&abs(<0.05"; TCut cTarget = "abs(<0.1"; TCut cCleanTrk = cGoodVDC&&cGoodAngles&&cTarget;

Cerenkov Efficiency Study

Prior to afternoon of 1 Nov 2010, Cerenkov threshold at 200 mV. Inefficiency = 7% After lowering the threshold to 50 mV, Cerenkov inefficiency < 1.5% (Updated 12 Nov 2010)



  • Current scan : (Runs: 1uA: 9250, 2uA: 9251, 4uA: 9252)
  • Buffered mode (10-28): [this halog]
  • info about livetime from TS registers [this halog]
  • Pulser runs for dead time studies

LHRS commissioning

Calibrate the ADCs:

Still to be done S1, Pion rejector, (Shloka is working on that)

[Cerenkov], [S2 (done 10-09)]

Calibrate detector time spectra:

Still to be done : understand timing spectra for S1,S2 and the Cerenkov

Evaluate PID efficiency:

LHRS optics run (no sieve/sieve)

Ye Zhihong is interested in this analysis and started it.

LHRS acceptance check

[delta scan on 10-22] need to be analyzed

DVCS Calorimeter

  • two photon invariant mass: [10-24]

Elastic calibration


BPM calibration

BCM calibration

  • partial calibration on December 23: [results], coda run number is 9875
  • partial calibration on December 13: [results]
  • third calibration on November 24: [epics calibration]
  • bcm gain jump (Nov 12): [halog]
  • second calibration on 11-05 (after temperature is fixed): [here]
  • the initial calibration was done on 10-14 [halog]. Calibration results are [here]

Beam energy measurement

  • the initial measurement was done on 10-15: [halog]



Analysis outputs

From Carlos:

You can find under the directory /adaql4/work1/dvcs/ntuples/ in the adaq computers ntuples corresponding to a first "fast" analysis pass.

Each file contains 2 ntuples (ntu1 and ntu2) corresponding to 1 and 2 clusters events. The name of the variables should be self-explanatory, but feel free to ask me.

These files correspond to a first analysis pass with a software calorimeter threshold of 1 GeV, fast waveform analysis (1-pulse fit only) and (elastic) calibration coefficients from Oct 26. Timing cut is -3,+3 ns wrt HRS. Additional cuts include 1-track only in the HRS and reasonable vertex reconstruction (within the target length).

All runs (even junk or LeftHRS-only, see PS below) are analyzed automatically as they are produced (although there is a backlog of a couple of days).

Please report any problems you might see, or any interesting variables missing.

Trigger efficiency studies

Run 9745 with singles